OpenStudy (anonymous):

Find the x intercept, identify the y intercept, and identify the slope. y=3/2x - 3

OpenStudy (tkhunny):

x-intercept corresponds to y = 0. Substitute and solve. y-intercept and slope can be read from the problem statement, since it is provided in teh Slope-Intercept form.

OpenStudy (anonymous):

So it's 0=3/2x-3 3=3/2x x=2

OpenStudy (anonymous):

And the slope is 2, but what about the y intercept?

OpenStudy (tkhunny):

x-intercept corresponds to y = 0. Substitute and solve. You found the x-intercept, (2,0) y=3/2x - 3 y-intercept (0,3) Slope: 3/2